what is the magnitude of current that flows across a 9-V battery connected to a resistor 15ohm? how long it will take a charge of 30 C to pass a cross section in a circuit? calculate the power rating.


1
Expert's answer
2020-07-16T11:14:59-0400

Given "V = 9V, R = 15\Omega"

Then by ohm's law,

current, "I = \frac{V}{R} = \frac{9}{15} = 0.6A"


Also, "I = \frac{Q}{t} \implies t = \frac{Q}{I}" "= \frac{30}{0.6} = 50 s"



Power, "P = VI = 9*0.6 = 5.4 W"